# TRACE Back from the Future: A Probabilistic Reasoning Approach to Controllable Language Generation

### v1

@@ Abstract
As large language models (LMs) advance, there is an increasing need to control their outputs to align with human values (e.g., detoxification) or desired attributes (e.g., personalization, topic). However, autoregressive models focus on next-token predictions and struggle with global properties that require looking ahead. Existing solutions either tune or post-train LMs for each new attribute—expensive and inflexible—or approximate the Expected Attribute Probability (EAP) of future sequences by sampling or training, which is slow and unreliable for rare attributes. We introduce TRACE (Tractable Probabilistic Reasoning for Adaptable Controllable gEneration), a novel framework that efficiently computes EAP and adapts to new attributes through tractable probabilistic reasoning and lightweight control. TRACE distills a Hidden Markov Model (HMM) from an LM and pairs it with a small classifier to estimate attribute probabilities, enabling exact EAP computation over the HMM’s predicted futures. This EAP is then used to reweigh the LM’s next-token probabilities for globally compliant continuations. Empirically, TRACE achieves state-of-the-art results in detoxification with only 10% decoding overhead, adapts to 76 low-resource personalized LLMs within seconds, and seamlessly extends to composite attributes.
@@ 1 Introduction
$p_{lm}=\textbf{0.3}$$\times$$=$$p_{\textit{TRACE}}\propto 0.03$to
@@
$p_{lm}=0.1$$\times$$=$$p_{\textit{TRACE}}\propto\textbf{0.08}$
@@
As large language models (LMs) become more ubiquitous in commercial products and daily life, there is a growing need to control their outputs. There is a large body of research in LM alignment with objectives such as detoxification (Gehman et al., 2020; Xu et al., 2021), where abundant data and benchmarks exist. Meanwhile, there is growing interest involving specialized or personal attributes under low-data conditions (Adiwardana et al., 2020; Xu et al., 2021), as well as compositional attributes for creating complex or novel outputs that rarely appear during standard training (Liu et al., 2023).
Controllable text generation is challenging because most language models are autoregressive, generating each token solely from its predecessors without looking ahead—yet many attributes depend on the entire text. One line of solutions modifies the base model’s distribution via fine-tuning or post-training (e.g., RL, RLHF) (Gururangan et al., 2020; Rafailov et al., 2024; Schulman et al., 2017), but these approaches can be highly expensive and risk degrading fluency or diversity (Kumar et al., 2022). Thus, there is a need for more lightweight solutions that can directly leverage pre-trained LMs by changing the decoding strategy.
@@
In this work, we propose a framework for controllable generation that utilizes tractable models to approximate the probability of satisfying an attribute given a partial sequence. This approach avoids costly sampling methods and eliminates the need for attribute-specific fine-tuning or retraining of the base language model (LM). First, we perform a one-time distillation of a tractable Hidden Markov Model (HMM) to approximate the base LM. Next, for each desired attribute, we train a log-linear classifier to estimate the probability of satisfying a target attribute given the entire text. During decoding, we use Bayesian conditioning to reweight next-token probabilities based on the likelihood that future sequences—predicted by the HMM—comply with these attributes, which we refer to as Expected Attribute Probability (EAP) (Figure 1). Crucially, the combination of the tractable HMM and log-linear classifier enables us to compute EAP efficiently and exactly over all of the exponentially many future sequences (over the HMM state space).
@@
We evaluate TRACE on three important tasks: (1) Detoxification: TRACE outperforms expensive RL, training- and sampling-based baselines on GPT2-large and Gemma 2B, taking only 1.1$\sim$ 1.1$\times$ 1.1 × time per-token to standard decoding. (2) Personalized LLMs: TRACE adapts to 76 distinct characters in about three seconds each, outperforming prompting approaches with only a few hundred training samples. (3) Compositional Attributes: We show how to easily generate conditional on compositions of attributes, which may be anti-correlated such as texts that are both political and nontoxic. Overall, TRACE is a simple, lightweight controllable generation approach that achieves state-of-the-art detoxification performance, extends to low-resource and composite text generation, and scales well to modern LLMs.
@@ 2 Related Work
Controllable text generation methods can be broadly categorized based on whether they modify the underlying language model through training or alter the decoding process of a pre-trained model.
@@ 2.1 Training Methods
One line of approaches modifies the base language model’s (LM) parameters, typically through fine-tuning or reinforcement learning, to instill desired attributes directly into the model’s distribution. DAPT (Gururangan et al., 2020) fine-tunes the base LM on domain-specific data. PPO (Schulman et al., 2017) uses a reward model and policy gradients to fine-tune the LM towards desired behaviors like non-toxicity, while DPO (Lee et al., 2024) aligns the model using pairwise preferences without an explicit reward model. Similarly, Quark (Lu et al., 2022) uses an RL-like procedure conditioned on a learned reward token. A major drawback of these methods is the need for substantial data and costly retraining of the large LM for each new attribute or set of attributes. Furthermore, modifying the base LM risks degrading its general fluency and diversity (Kumar et al., 2022).
@@ 2.2 Decoding Methods
As directly computing the EAP requires summing over exponentially many future sequences, one approach trains auxiliary models (discriminators or guides) to predict the likelihood of satisfying an attribute based only on the partial sequence generated so far. FUDGE (Yang & Klein, 2021) trains a discriminator that predicts EAP to modify next-token probabilities. DExperts (Liu et al., 2021) combines pre-trained ”expert” and ”anti-expert” LMs, while GeDi (Krause et al., 2021) trains smaller attribute-conditional LMs (guides) to reweigh the base LM’s distribution via Bayes’ rule. LiSeCo (Cheng et al., 2024) uses a learned linear probe in latent space. A fundamental challenge here is that accurately predicting the final attribute satisfaction from a prefix requires the auxiliary model to implicitly reason about likely continuations and their properties, effectively needing sophisticated lookahead capabilities that are difficult to train reliably. Furthermore, these methods still typically require training a separate, often large, auxiliary model for each attribute.
@@
While all methods attempting EAP-based control must approximate the true (intractable) EAP under the base LM, the nature of the approximation differs significantly. Discriminator/guide-based and sampling-based methods directly approximate the EAP value during decoding, often requiring complex per-attribute training or incurring high variance and computational cost at generation time. In contrast, TRACE shifts the primary approximation to the one-time HMM distillation step ($p_{\text{hmm}}\approx p_{\text{lm}}$). Conditioned on this distilled HMM, TRACE then performs an exact and tractable computation of the EAP ($p_{\text{hmm}}(s\mid x_{<t},x_{t})$) relative to the HMM distribution, as detailed in Section 4.2. This design choice—approximating the LM via a one-time HMM distillation—enables efficient, adaptable, and low-variance EAP calculation during decoding, avoiding per-attribute retraining or costly sampling.
@@ 2.3 Control via Tractable Models
Tractable probabilistic models (Choi et al., 2020) such as HMMs enable efficient computation of various quantities such as marginals and the probability of satisfying logical constraints, while such computations on autoregressive models are often provably hard even to approximate (Roth, 1996). They have been effectively used, for instance, to enforce adherence to logical or lexical requirements in generative modeling (Liu et al., 2024a, c). In the context of language modeling, Ahmed et al. (2023) proposed using a local tractable approximation for training autoregressive models to satisfy logical constraints. Meanwhile, Zhang et al. (2023, 2024) proposed utilizing HMM structures specifically to enforce logical constraints such as those given by deterministic finite automata (DFA).
While powerful for such formally specifiable constraints, this approach does not readily extend to controlling generation based on high-level semantic properties like style, safety, or persona, which lack simple logical definitions and depend on the overall meaning of the text. Our work, TRACE, bridges this gap by employing the HMM uniquely for semantic control, enabling efficient probabilistic reasoning about semantic attributes (via EAP computation) rather than enforcing logical rules.
@@ 3.1 Controllable Generation
We consider generating a text (sequence of tokens) $x_{1:n}$ of length $n$ from a large language model (LM). In controllable generation, the goal is to generate text from the LM conditional on some attribute $s$, such as nontoxicity. We assume that the attribute can be measured by some probabilistic classifier $p(s|x_{1:n})\in[0,1]$ representing the probability (or degree) of satisfaction of the attribute given the full text.
@@ 4.1 TRACE: Guiding LM with HMM Probabilities
Here, $p_{\textit{hmm}}(s\mid x_{<t},x_{t})$ is the probability, under our HMM and attribute classifier, that the entire future continuation will satisfy the attribute. In contrast to the formulation of Zhang et al. (2023), $s$ is not a logical attribute that maps to 00 or $1$, but instead represents a semantic attribute given by a probabilistic classifier $p(s|x_{1:n})$. Clearly, if the classifier $p(s|x_{1:n})$ is arbitrary without any structure (e.g., a neural network), then computing the expected attribute probability will be intractable as we will again need to generate all possible continuations to feed to the classifier.
@@ 4.2 Tractable Computation of EAP
$=\sum_{z_{t}}p_{\textit{hmm}}(z_{t}\mid x_{\leq t})\boxed{\sum_{x_{>t}}p_{\textit{hmm}}(x_{>t}|z_{t})\cdot p(s|x_{>t},x_{\leq t})}.$
@@
with the base case $p_{\textit{hmm}}(z_{1},x_{\leq 1})=p(z_{1})p(x_{1}|z_{1})$. To obtain the conditional, we simply divide by the normalizing constant $p_{\textit{hmm}}(x_{\leq t})=\sum_{z_{t}}p_{\textit{hmm}}(z_{t},x_{\leq t})$ Note that this computation is independent of the classifier.
@@
Integration During Generation Algorithmically, as each new token is generated, the forward probability $p(z_{t}\mid x_{<t},x_{t})$ is updated based on the recursion in Equation 6. Using the precomputed backward expectations, the overall expected attribute probability is computed as
$p_{\textit{hmm}}(s\mid x_{t},x_{<t})$
$=\left(\prod_{i\leq t}w(x_{i})\right)\sum_{z_{t}}p_{\textit{hmm}}(z_{t}|x_{\leq t})\cdot\mathbb{E}_{\textit{hmm}}\left[\prod_{i>t}w(x_{i})\Bigm{|}z_{t}\right].$ (8)
Plugging this into Equation 5, we can then compute the constrained next-token distribution:
$p_{\text{TRACE}}(x_{t}\mid x_{<t},s)\propto p_{\textit{lm}}(x_{t}\mid x_{<t})\cdot p_{\textit{hmm}}(s\mid x_{t},x_{<t})$ (9)
@@
Table 1: Detoxification performance on RealToxicityPrompts (Gehman et al., 2020), using 10k prompts for GPT-2 and 1k for Gemma-2B. TRACE uses training-time transformation only; +Dec. TF adds decoding-time transformation (Section 4.3); HMM uses an HMM trained on 1M vs. 10M samples (Appendix. B). Struck-through fluency often signifies unnatural repetition (Holtzman et al., 2020). Baselines: (1) Gururangan et al. (2020); (2) Krause et al. (2021); (3) Yang & Klein (2021); (4) Liu et al. (2021); (5) Dathathri et al. (2020); (6) Kumar et al. (2022); (7) Schulman et al. (2017); (8) Lu et al. (2022); (9) Lee et al. (2024) (Implemented by us, see Appendix D).
Model Toxicity (↓) Fluency (↓) Diversity (↑) Approach Type
@@
GPT2 0.385 0.254 25.57 0.87 0.86 Baseline
DAPT(1) 0.428 0.360 31.21 0.84 0.84 Finetuning
GeDi(2) 0.363 0.217 60.03 0.84 0.83 Decoding (Trained Guide)
FUDGE(3) 0.302 0.371 12.97* 0.78 0.82 Decoding (Trained Guide)
DExperts(4) 0.314 0.128 32.41 0.84 0.84 Decoding (Trained Guide)
PPLM(5) 0.520 0.518 32.58 0.86 0.86 Decoding (Logit Control)
MuCoLa(6) 0.308 0.088 29.92 0.82 0.83 Decoding (Sampling)
PPO(7) 0.218 0.044 14.27* 0.80 0.84 RL
Quark(8) 0.196 0.035 12.47* 0.80 0.84 RL
DPO(9) 0.180 0.026 21.59* 0.76 0.78 RL
TRACE 0.187 0.026 27.51 0.87 0.85 Decoding (HMM Reasoning)
TRACE (+Dec. TF) 0.163 0.016 29.83 0.85 0.85 Decoding (HMM Reasoning)
@@
Gemma-2B 0.359 0.23 15.75 0.86 0.85 Baseline
DPO(9) 0.222 0.06 14.39* 0.74 0.77 RL
TRACE (HMM) 0.195 0.03 16.78 0.86 0.85 Decoding (HMM Reasoning)
TRACE 0.189 0.02 17.68 0.86 0.85 Decoding (HMM Reasoning)
4.3 Attribute Classifier Fitting
As discussed, one class of classifiers that make EAP computation tractable is a product over tokens, that is linear in log space. We show that attributes are not fully factorizable to varying extents (See 3). Nonetheless, we demonstrate later in Section 5.2 that even such a simple classifier, when combined with a HMM, captures EAP sufficiently accurately to outperform even baselines using neural classifiers (e.g., GeDi, FUDGE).
To fit such classifiers, we assume the availability of an oracle $p_{\text{oracle}}$ that scores the likelihood of any text $x_{1:n}$ satisfying a given attribute $s$. This oracle can be a classifier trained on human-labeled data or an existing API-based measure of attribute satisfaction.
Fitting Rare attributes via Log-MSE We target attributes that may be rare in large text corpora (e.g., toxic or political content). To effectively model these rare attributes, we define attributes as the absence of these properties (e.g., nontoxicity). Then, we aim to penalize mispredictions of near-zero scores (e.g., toxic texts) more heavily. To do this, instead of using the standard cross-entropy loss, we minimize a mean squared error loss in log-space (log-MSE), which amplifies error penalties in the low-probability regions. Specifically, for our factorized (log-linear) classifier, the log-probability of the attribute $s$ given the text $x_{1:n}$ is a sum of log-weights, $\log p(s\mid x_{1:n})=\sum_{i=1}^{n}\log w(x_{i})$. Then, given oracle scores $p_{\text{oracle}}$ (potentially transformed, as discussed next), the loss for a text $x$ is given by
Transforming Oracle Probabilities In practice, it is often beneficial to transform the raw oracle probability scores $p_{\text{oracle}}$ to enforce stricter attribute satisfaction and achieve a clearer separation between acceptable and unacceptable outputs. For example, if an oracle assigns a text a 20% probability of being non-toxic, we might transform this score to a lower probability, encouraging safer generation.
We apply an affine transformation in the logit space using non-negative scalars $b$ (scale) and $c$ (shift). The transformed probability $p^{\prime}_{\text{oracle}}$ is computed as
@@ 4.3 Attribute Classifier Fitting
where $\sigma(z)=1/(1+e^{-z})$ is the standard sigmoid function. This transformation allows fine-grained control over the probability distribution’s shape. As discussed in Section 5.2, applying this transformation enhances the model’s ability to distinguish between attribute classes and improves downstream control during text generation.

### v2

@@ Abstract
As large language models (LMs) advance, there is an increasing need to control their outputs to align with human values (e.g., detoxification) or desired attributes (e.g., personalization, topic). However, autoregressive models focus on next-token predictions and struggle with global properties that require looking ahead. Existing solutions either post-train LMs for each new attribute—expensive and inflexible—or approximate the Expected Attribute Probability (EAP) of future sequences by sampling or training, which is slow and unreliable for rare attributes. We introduce TRACE (Tractable Probabilistic Reasoning for Adaptable Controllable gEneration) [^1] Our code is available at: https://github.com/yidouweng/trace, a novel framework that efficiently computes EAP and adapts to new attributes through tractable probabilistic reasoning and lightweight control. TRACE distills a Hidden Markov Model (HMM) from an LM and pairs it with a small classifier to estimate attribute probabilities, enabling exact EAP computation over the HMM’s predicted futures. This EAP is then used to reweigh the LM’s next-token probabilities for globally compliant continuations. Empirically, TRACE achieves state-of-the-art detoxification results with only 20% decoding overhead, yields 76 low-resource personalized LMs within seconds, and seamlessly extends to composite attributes.
@@ 1 Introduction
$p_{lm}=\textbf{0.3}$$\times$$p_{\textit{TRACE}}\propto 0.03$to
@@
$p_{lm}=0.1$$\times$$p_{\textit{TRACE}}\propto\textbf{0.08}$
@@
As large language models (LMs) become more ubiquitous in commercial products and daily life, there is a growing need to control their outputs. There is a large body of research in LM alignment with objectives such as detoxification (Gehman et al., 2020; Xu et al., 2021), where abundant data and benchmarks exist. Meanwhile, there is growing interest involving specialized or personal attributes under low-data conditions (Adiwardana et al., 2020; Xu et al., 2021), as well as compositional attributes for creating complex or novel outputs that rarely appear during standard training (Liu et al., 2022).
Controllable text generation is challenging because most language models are autoregressive, generating each token solely from its predecessors without looking ahead—yet many attributes depend on the entire text. One line of solutions modifies the base model’s distribution via fine-tuning or post-training (e.g., RL, RLHF) (Gururangan et al., 2020; Rafailov et al., 2023; Schulman et al., 2017), but these approaches can be highly expensive and risk degrading fluency or diversity (Kumar et al., 2022). Thus, there is a need for more lightweight solutions that can directly leverage pre-trained LMs by changing the decoding strategy.
@@
In this work, we propose a framework for controllable generation that utilizes tractable models to approximate the probability of satisfying an attribute given a partial sequence. This approach avoids costly sampling methods and retraining the base language model (LM) for each target attribute. First, we perform a one-time distillation of a tractable Hidden Markov Model (HMM) to approximate the base LM. Next, for each desired attribute, we train a log-linear classifier to estimate the probability of satisfying a target attribute given the entire text. During decoding, we use Bayesian conditioning to reweight next-token probabilities based on the likelihood that future sequences—predicted by the HMM—comply with these attributes, which we refer to as Expected Attribute Probability (EAP) (Figure 1). Crucially, the combination of the tractable HMM and log-linear classifier enables us to compute EAP over all of the exponentially many future sequences efficiently and exactly (over the HMM state space).
@@
We evaluate TRACE on three important tasks: (1) Detoxification: TRACE outperforms expensive RL, training- and sampling-based baselines on GPT2-large and Gemma 2B. (2) Personalized LLMs: TRACE adapts to 76 distinct characters in about three seconds each, outperforming prompting approaches with only a few hundred training samples, and taking only $\sim 1.2\times$ time per-token to standard decoding. (3) Compositional Attributes: TRACE seamlessly generates texts satisfying multiple attributes–e.g. political and nontoxic–a combination too sparse for training or sampling methods. Overall, TRACE is a simple, lightweight controllable generation approach that achieves state-of-the-art detoxification performance, extends to low-resource and composite text generation, and scales well to modern LLMs.
@@ 2 Related Work
Controllable text generation methods fall into two main categories: those that modify the language model (LM) via training, and those that steer a pre-trained model through decoding-time interventions.
@@ 2.1 Training Methods
One line of approaches modifies the base LM parameters, typically through fine-tuning or reinforcement learning, to instill desired attributes directly into the model’s distribution. DAPT (Gururangan et al., 2020) fine-tunes the base LM on domain-specific data. PPO (Schulman et al., 2017) uses a reward model and policy gradients to fine-tune the LM towards desired behaviors like non-toxicity, while DPO (Lee et al., 2024) aligns the model using pairwise preferences without an explicit reward model. Similarly, Quark (Lu et al., 2022) uses an RL-like procedure conditioned on a learned reward token. A major drawback of these methods is the need for substantial data and costly retraining of the large LM for each new attribute or set of attributes. Furthermore, modifying the base LM risks degrading its general fluency and diversity (Kumar et al., 2022).
@@ 2.2 Decoding Methods
Since exact EAP computation requires summing over future sequences, many methods train auxiliary models to estimate attribute satisfaction from partial generations. FUDGE (Yang & Klein, 2021) and NADO (Meng et al., 2022) train discriminators to approximate EAP; DExperts (Liu et al., 2021) blends expert and anti-expert LMs; GeDi (Krause et al., 2021) uses attribute-conditioned guides with Bayes’ rule; and LiSeCo (Cheng et al., 2024) applies a linear probe in latent space. A key challenge across these methods is that accurately estimating future attribute satisfaction from prefixes requires nontrivial lookahead, and training a separate large, auxiliary model per attribute adds significant overhead.
@@
[…]
@@ 2.3 Control via Tractable Models
Tractable probabilistic models (Choi et al., 2020) such as HMMs enable efficient computation of various quantities such as marginals and the probability of satisfying logical constraints–computations that are provably hard even to approximate on autoregressive models (Roth, 1996). Prior work has used such models to enforce logical or lexical constraints in generative modeling (Liu et al., 2024a, b). In language modeling, Ahmed et al. (2023) proposed using local tractable approximations for training autoregressive models to satisfy logical constraints. Meanwhile, Zhang et al. (2023, 2024) used HMMs to enforce logical constraints such as those given by deterministic finite automata (DFA).
While powerful for such formally specifiable constraints, this approach does not readily extend to high-level semantic attributes such as style, safety, or persona, which lack symbolic definitions and depend on the overall meaning of the text. Our work, TRACE, bridges this gap by employing the HMM uniquely for semantic control, enabling efficient probabilistic reasoning about semantic attributes (via EAP computation) rather than enforcing logical rules.
@@ 3.1 Controllable Generation
We consider generating a text (sequence of tokens) $x_{1:n}$ of length $n$ from a language model (LM). In controllable generation, the goal is to generate text from the LM conditional on some attribute $s$, such as nontoxicity. We assume that the attribute can be measured by some probabilistic classifier $p(s|x_{1:n})\in[0,1]$ representing the probability (or degree) of satisfaction of the attribute given the full text.
@@ 4.1 TRACE: Guiding LM with HMM Probabilities
Here, $p_{\textit{hmm}}(s\mid x_{<t},x_{t})$ is the probability, under our HMM and attribute classifier, that the entire future continuation will satisfy the attribute. In contrast to the formulation of Zhang et al. (2023), $s$ is not a logical attribute that maps to 0 or $1$, but instead represents a semantic attribute given by a probabilistic classifier $p(s|x_{1:n})$. Clearly, if the classifier $p(s|x_{1:n})$ is arbitrary without any structure (e.g., a neural network), then computing the expected attribute probability will be intractable as we will again need to generate all possible continuations to feed to the classifier.
@@ 4.2 Tractable Computation of EAP
$=\sum_{z_{t}}p_{\textit{hmm}}(z_{t}\mid x_{\leq t})\boxed{\sum_{x_{>t}}p_{\textit{hmm}}(x_{>t}\mid z_{t})\,p(s\mid x_{>t},x_{\leq t})}.$ (6)
@@
with the base case $p_{\textit{hmm}}(z_{1},x_{\leq 1})=p(z_{1})p(x_{1}|z_{1})$. To obtain the conditional, we simply divide by the normalizing constant $p_{\textit{hmm}}(x_{\leq t})=\sum_{z_{t}}p_{\textit{hmm}}(z_{t},x_{\leq t})$ . Note that this computation is independent of the classifier.
@@
4.3 Attribute Classifier Fitting
As discussed, one class of classifiers that make EAP computation tractable is a product over tokens, that is, linear in log space. In practice, attributes are not fully factorizable (to varying extents); nonetheless, we show that even such a simple classifier, when combined with a HMM, captures EAP sufficiently accurately to outperform baselines using neural classifiers (e.g., GeDi, FUDGE) (See 5.2).
We assume access to an oracle $p_{\text{oracle}}$ that scores the probability of a sequence $x_{1:n}$ satisfying a target attribute $s$. This oracle may be a trained classifier (e.g., from human annotations) or an external API scoring attribute satisfaction.
Fitting Rare attributes via Log-MSE Our goal is to model attributes that are rare in natural text, such as toxicity or political content. To effectively identify these cases, the learning objective must distinguish texts that exhibit the rare attribute from the more common, neutral ones. Standard objectives like cross-entropy treat all misclassifications symmetrically, which is suboptimal when one class is rare.
We therefore use an asymmetric loss that more heavily penalizes misclassifying the rare examples. In particular, we use mean squared error loss in log-space (log-MSE), which amplifies error penalties in the low-probability region. To align with this objective, we define each attribute by its absence (e.g., modeling ”nontoxicity” rather than ”toxicity”).
Specifically, for our factorized (log-linear) classifier, the log-probability of the attribute $s$ given the text $x_{1:n}$ is a sum of log-weights, $\log p(s\mid x_{1:n})=\sum_{i=1}^{n}\log w(x_{i})$. Then, given oracle scores $p_{\text{oracle}}$ (potentially transformed, as discussed next), the loss for a text $x$ is given by
Table 1: Detoxification on RealToxicityPrompts (Gehman et al., 2020). Results are shown for GPT-2 (10k prompts) and Gemma-2B (1k prompts). TRACE applies training- and decoding-time transformation (see Section 4.3). Struck-through fluency indicates unnatural repetition (Holtzman et al., 2020). Baselines: (1) Gururangan et al. (2020); (2) Krause et al. (2021); (3) Yang & Klein (2021); (4) Liu et al. (2021); (5) Dathathri et al. (2020); (6) Kumar et al. (2022); (7) Schulman et al. (2017); (8) Lu et al. (2022); (9) Lee et al. (2024) (our implementation).
Model Toxicity (↓) Diversity (↑) PPL (↓) Approach Type
@@ 4.3 Attribute Classifier Fitting
GPT2 0.385 0.254 0.87 0.86 25.57 Baseline
DAPT(1) 0.428 0.360 0.84 0.84 31.21 Finetuning
GeDi(2) 0.363 0.217 0.84 0.83 60.03 Decoding (Trained Guide)
FUDGE(3) 0.302 0.371 0.78 0.82 12.97* Decoding (Trained Guide)
DExperts(4) 0.314 0.128 0.84 0.84 32.41 Decoding (Trained Guide)
PPLM(5) 0.520 0.518 0.86 0.86 32.58 Decoding (Logit Control)
MuCoLa(6) 0.308 0.088 0.82 0.83 29.92 Decoding (Sampling)
PPO(7) 0.218 0.044 0.80 0.84 14.27* RL
Quark(8) 0.196 0.035 0.80 0.84 12.47* RL
DPO(9) 0.180 0.026 0.76 0.78 21.59* RL
TRACE 0.163 0.016 0.85 0.85 29.83 Decoding (HMM Reasoning)
@@
Gemma-2B 0.359 0.23 0.86 0.85 15.75 Baseline
DPO(9) 0.222 0.06 0.74 0.77 14.39* RL
TRACE 0.189 0.02 0.86 0.85 17.68 Decoding (HMM Reasoning)
Probability Transformation for Better Control
In practice, it can be beneficial to transform raw oracle probability scores ($p_{\text{oracle}}$) for a clearer separation between desired and undesired outputs, and enforce stricter attribute satisfaction. For instance, a 20% non-toxicity score from an oracle can be transformed to a lower value to encourage safer generation.
To do this, we propose to apply an affine transformation to the oracle scores in logit space using a (non-negative) scale ($b$) and shift ($c$):
@@
where $\sigma(z)$ is the sigmoid function. This transformation reshapes the target distribution; specifically, increasing the scale $b$ pushes intermediate probabilities towards the extremes of 0 and 1. This creates a more bimodal distribution, which helps to clearly distinguish between attribute-compliant and non-compliant texts.
This probability transformation can also be applied at decoding time to the EAP, $p_{\text{hmm}}(s\mid x_{t},x_{<t})$, before it influences token generation (Eq. 5). This can be used to sharpen the separation between next tokens $x_{t}$ based on their EAP, thus ensuring stricter control of the attribute.
In Section 5.2, we empirically justify this intuition and carefully study the effect of these training-time and decoding-time probability transformations.